OpenStudy (anonymous):

write the trig expression in terms of sine and cosine then simplify. cos t tan t cos t csc t

OpenStudy (accessdenied):

\[\tan t = \frac{\sin t}{\cos t} \\ \cot t = \frac{\cos t}{\sin t} \\ \sec t = \frac{1}{\cos t} \\ \csc t = \frac{1}{\sin t} \] Just some definitions that would help. You basically just substitute the definition with sines/cosines and simplify from there.

OpenStudy (anonymous):

Ok so what's the answer?

OpenStudy (accessdenied):

Try doing the problem as I said and find out. ;)

OpenStudy (anonymous):

im so lost

OpenStudy (accessdenied):

We have: \(\cos t \tan t\) By definition, \(\large{ \tan t = \frac{\sin t}{\cos t} } \) So, we can substitute the definition of \(\tan t\) for it in the expression: \[ \cos t \times \frac{\sin t}{\cos t} = \frac{\sin t \cos t}{\cos t}\] There, you can simplify further because the \(\cos t\)'s will have both one in the numerator and denominator.